TOWARDS AN ACCURATE INFRARED LINELIST FOR CO_{2} AND ISOTOPOLOGUES

\underline{TIMOTHY J. LEE}, MS 245-1, NASA Ames Research Center, Moffett Field, CA, 94035; XINCHUAN HUANG, SETI Institute, 189 Bernardo Ave, Suite 100, Mountain View, CA, 94043; DAVID W. SCHWENKE, MS T27B-1, NASA Ames Research Center, Moffett Field, CA, 94035; and SERGEY TASHKUN, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics, SB, Russian Academy of Science, 634055, Tomsk, Russia.

Following the ”Best Theory + High-resolution Experimental Data” strategy, we have completed an initial CO_{2} infrared (IR) line list, denoted Ames-1. A procedure similar to the one used for ammonia^{a} is adopted to generate a global potential energy surface (PES), including various small corrections such as scalar relativity, extrapolation to the one-particle basisset limit, and a higher-order correlation correction, followed by refinement using accurate high-resolution laboratory data. We will discuss limitations in the use of HITRAN data for the refinement step. The current PES yields uncertainties of 0.01-0.02 cm^{−1}for J =0-117 ^{12}C^{16}O_{2} energy levels up to 13,845 cm^{−1} above the zero-point level. Comparisons between Ames-1, HITRAN, and CDSD will be presented for room temperature as well as selected higher temperatures. Results for 13C, ^{14}C, ^{17}O, and ^{18}O isotopologues will be presented, and the accuracy for isotopologues will be discussed. Limitations of the current PES and the Ames-1 line list together with strategies to improve these will be discussed.

aX. Huang, D.W. Schwenke, and T.J. Lee, J. Chem. Phys. \underline{129}, 214304 (2008); J. Chem. Phys. \underline{134}, 044320/044321 (2011).